OpenStudy (anonymous):

How many milliliters of sodium metal, with a density of 0.97 g/mL, would be needed to produce 43.6 grams of sodium hydroxide in the single replacement reaction below? Be sure to show the work that you did to solve this problem.

OpenStudy (anonymous):

unbalanced equation: Na + H2O yields NaOH + H2

OpenStudy (anonymous):

is my working correct

OpenStudy (anonymous):

balanced equation: 2Na + 2H2O yields 2NaOH + H2 43.6gNaOH*1molNaOH/40gNaOH=1.09molNaOH*2molNa/2molNaOH=1.09molNa*23gNa/1molNa=25.0gNa density=mass/volume volume=mass/density volume Na=25g/0.97g/mL=25.8mL

OpenStudy (anonymous):

2Na + 2H2O ------------ 2NaOH + H2 molar mass 46 2 mass ? 43.6 then the mass of sodium=46×43.6÷2=1002.8 g Then the volume=1002.8/0.97=1033.8 ml